OpenStudy (anonymous):

What is the surface area of a sphere with radius of 3.5? A. 12.25p B. 14p C. 24.5p D. 49p

jimthompson5910 (jim_thompson5910):

use the formula SA = 4*pi*r^2 where SA = surface area r = radius

OpenStudy (anonymous):

i did and got 153.86

jimthompson5910 (jim_thompson5910):

leave pi as it is

OpenStudy (anonymous):

A spherical water tank has a diameter of 10 feet. How much water does it hold when half full? (Use 3.14 for p)

jimthompson5910 (jim_thompson5910):

find the volume of the full tank using V = (4/3)*pi*r^3 then cut that volume in half (since we want to know how much water there is when it's half full)